# SP-Chain: Boosting Intra-Shard and Cross-Shard Security and Performance in Blockchain Sharding

### v1

@@ 6.1 Epoch Security
6.2 Analysis of Randomness Generation
In this part, we mainly analyze our chain-based distributed randomness generation scheme’s security and performance, as it affects the leader rotation results. We will show that our scheme possesses unbiasability, unpredictability, verifiability, and scalability [32].
Unbiasability: The generated randomness represents an unbiased, uniformly random value, except with negligible probability. In our chain-based distributed randomness generation scheme, the slot information $t$ is publicly known and cannot be manipulated. The random oracle that uniformly maps the seed to the leader is also publicly known, so it cannot be manipulated and corrupted by a single node. Therefore, we will show that the signature information used in the randomness generation process cannot be biased, except with negligible probability.
First, according to Sec. 3, each newly joined node generates a Sybil-resistant identity, and is assigned a key pair by PKI, and will be randomly assigned to a specific shard. Second, according to the shard reconfiguration, nodes will update their shard member table before each epoch. Therefore, except with negligible probability, a leader cannot generate multiple legal identities/signatures in the same shard. Even if a malicious leader generates multiple legal signatures in some ways (e.g., bribing other node’s identity), as long as the total number of malicious nodes in a single shard does not exceed $k/2$, our system can guarantee security. In this case, the randomness generation process still cannot be biased. Because a malicious node cannot control the signature output, and the signature needs to be verified by all nodes in the shard. Therefore, the chain-based distributed randomness generation scheme is unbiased.
Unpredictability. No party learns anything about the generated randomness, except with negligible probability, until the last confirmed block is determined. At the beginning of slot $t$, each node will use the signature information on the latest confirmed block to determine the new leader. Therefore, the new randomness cannot be predicted until the last confirmed block is determined. The decision of the last confirmed block is carried out in $t-1$ slot. Therefore, before the start of $t-1$ slot, the new randomness cannot be predicted except with negligible probability.
Verifiability. The generated randomness is third-party verifiable. The random oracle, shard member information, leader candidate list, and the generated blocks are all publicly known. Therefore, anyone can verify the randomness generation process.
Scalability. The randomness generation process is highly scalable. Since our random number generation scheme only requires nodes to obtain information from the past blockchain and perform distributed calculations, no additional communication overhead is introduced during the randomness generation process. This is a considerable advantage compared to the $O(k^{2})$ communication overhead introduced in the previous works. Therefore, our randomness generation mechanism is highly scalable.
6.3 Analysis of Intra-Shard Consensus
We now analyze the security of our intra-shard consensus, i.e., the two-phase concurrent voting protocol and the leader rotation scheme.
Safety of Intra-Shard Consensus. Safety here means that all honest nodes in each shard agree on the same state of the shard (i.e., no forks or inconsistencies in each shard). We will prove that any honest node can receive the votes from all other honest nodes within the waiting time $\triangle_{d}+\triangle_{b}+\triangle_{v}$, and reach a consensus. Similar to the previous assumptions, our intra-shard consensus is based on the synchronous network model. Suppose that an honest node $i_{first}$ receives the digest first at time $t_{0}$, and another honest node $i_{last}$ receives the digest last at time $t_{0}+\delta$. It is obvious that $\delta\leq\triangle_{d}$. Suppose that in the worst case, $\delta=\triangle_{d}$. Then when $i_{last}$ waits for $\triangle_{b}$ and starts voting, $i_{first}$ has waited for $\triangle_{b}+\triangle_{d}$ (has started voting and broadcasted its vote). At this time, $i_{first}$ needs to wait for $\triangle_{v}$ to ensure that it can receive the vote broadcasted by $i_{last}$. Therefore, it can be proved that any honest node can receive the votes of all honest nodes and reach a consensus during the waiting time.
Our protocol is secure under various malicious behaviours. First, when a malicious leader tries to send messages with different values to different nodes (equivocation attack), all the honest nodes will notice that via different block digests. If the equivocation attack is detected, each node immediately launches the voting process without waiting for $\triangle_{b}$, and the voting procedure can still be finished within the waiting time. When a malicious node tries to send different votes, this behavior will also be discovered by honest nodes and the malicious node’s vote will be treated as invalid. Second, when a malicious leader fails to send messages (silence attack), then none of the nodes in the shard will receive the valid block during the waiting time. After the waiting time (i.e., period of one slot), a new leader will be elected and a new round of consensus will automatically start. When malicious nodes launch silence attacks, this behavior will not affect consensus reaching as long as the number of per-shard malicious nodes does not exceed $k/2$. This is because honest nodes can always collect enough valid votes to reach consensus within the waiting time. Third, when a malicious leader (or node) tries to send invalid messages (e.g., transaction manipulation, forging signatures), honest nodes will detect such malicious behaviours during voting process, since nodes will verify every received messages. In summary, when the malicious nodes in the shard does not exceed $k/2$, the malicious behaviours cannot pass our intra-shard consensus, our two-phase concurrent voting thus maintain safety against various typical Byzantine behaviours.
Liveness of Intra-Shard Consensus. Liveness here represents the system’s responsiveness and that it keeps making progress (i.e., all valid blocks are eventually added to the chain of each shard). Note that the expectation round for an honest leader is around two in the worst case ($<k/2$ malicious nodes per shard). According to the leader rotation scheme we proposed, no matter the current leader is good or bad, a new leader will be automatically elected at the beginning of a new slot. In our intra-shard consensus protocol, every honest node has the same view on the validity of the proposed block in each slot. Therefore, no matter whether the leader is malicious or not, all honest nodes will automatically elect the next leader after the voting period. Therefore, a valid block will eventually be produced by an honest leader.
Network condition may change over time. SP-Chain applies similar solutions to previous works [37, 13] to alleviate such responsiveness issue of synchronous consensus. Specifically, SP-Chain runs a pre-scheduled consensus among shard members about every week to agree on a new $\Delta$ so that the system adjusts its consensus speed with the latest average delay of the network. It can make the protocol responsive to long-term, more robust changes of the network as technology advances, ensuring that messages will be delivered to the whole shard within the time limit. If the network changes unexpectedly, nodes may find that the intra-shard consensus is temporarily suspended, as they may not receive enough votes during the waiting time. They will then agree on a new $\Delta$ to recover the protocol in time. As a result, our intra-shard consensus can alleviate the responsiveness issue and ensure liveness.
6.4 Analysis of Cross-Shard Transactions
We now analyze the security and overhead of our cross-shard transaction processing scheme.
Security of Cross-Shard Transactions. In SP-Chain, cross-shard transactions can guarantee eventual atomicity like other works [33, 1]. Similar to the design of [33], cross-shard transactions are divided into two parts: withdrawal and deposit. In the source shard, the sender’s fund will be withdrawn first. When the source shard confirms the withdrawal, the deposit part will be sent to the destination shard and wait for execution. It is possible that the destination shard’s current leader is malicious and does not execute the deposit. However, sooner or later, an honest leader will be elected to execute the deposit and eventually achieve the transaction’s atomicity. Another possibility is that the leader in the source shard does not send the deposit part. In this case, the clients (users who send transactions) can inform the shard node that it has not received the corresponding deposit. So that the deposit part is resent, and the eventual atomicity is achieved.
In the process of cross-shard transactions, malicious leaders may tamper with cross-shard transactions. In this case, the destination shard nodes can verify whether the cross-shard transactions have been tampered with through the Merkle proof. Moreover, in our scheme, each node updates and maintains the shard member table of all shards in each epoch. Therefore, the malicious leader cannot deceive the destination shard’s nodes by forging the signature information. In summary, our cross-shard transaction processing scheme guarantees security.
@@ 6.4 Analysis of Cross-Shard Transactions
7 Implementation and Evaluation
7.1 Experimental Setup
We implement SP-Chain based on Harmony [1], one of the most advanced and well-known public blockchain sharding projects within top 50 market cap in cryptocurrency. SP-Chain is implemented in Go language with 5,000+ lines of code. We implement BLS aggregated signature [3] in the prototype to reduce the signature size for better performance. As a public blockchain sharding system, we implement a similar incentive mechanism to Harmony in our prototype. We choose to mainly use Harmony as the baseline protocol to compare the performance with SP-Chain. The main protocols in SP-Chain can be easily applied to most existing sharding systems to improve system performance. We choose Harmony as the baseline mainly for fair comparison. Besides Harmony, we also make a generic comparison of SP-Chain with some other related works.
The choice of the experimental environment is similar to that in previous work [37]. Specifically, we deploy SP-Chain on Amazon EC2 with up to 32 machines, each running up to 125 SP-Chain nodes. Therefore, the total network size scales up to 4,000 nodes. The machine is selected as c5.24xlarge, each with a 96-core processor and a 25-Gbps communication link. To simulate geographically-distributed nodes, we consider a latency of 100 ms for every message and a bandwidth of 20 Mbps for each node. In each shard, we assume that when the adversaries observe the leader, they have the ability to attack it in the next slot. Moreover, in each shard, the total malicious nodes are less than half of all nodes in a shard. We set each transaction size to 512 bytes, and each block contains up to 4,096 transactions, resulting in a block size of 2MB.
@@ 7.1 Experimental Setup
Choice of Shard Size. We first determine the number of nodes per shard based on Equation 2. The rule for selecting the shard size is: the failure probability of each shard should be less than $2^{-20}\approx 9\cdot 10^{-7}$ (time-to-failure of more than 4580 years for one-day epoch) [9, 37]. Table I shows the choice of shard size under different shard numbers and the corresponding failure probability. Results show that our choice of shard size makes the probability of failure less than $9\cdot 10^{-7}$ at any scale, ensuring the security of the system. The following experiments will be conducted based on the shard size determined by Table I.
@@
Choice of $\triangle$. The broadcast upper bound latency of digest, block and vote ($\triangle_{d}$, $\triangle_{b}$ and $\triangle_{v}$) are the 3 most important parameters in our system. To determine them, we evaluate the actual time spent in broadcasting digests, blocks, and votes in the system, and show the experimental results in Fig. 4. Specifically, the bottom/top of each bar is the minimum/maximum delay in the broadcast, and the horizontal line connected to each bar is the average delay. Based on the measured actual broadcast delay, we set the broadcast delay upper bound as $delay_{max}+\sigma$. Where $delay_{max}$ is the actual maximum broadcast delay, $\sigma$ is the standard deviation obtained by fitting the broadcast delay to a normal distribution. Table II shows the calculated delay upper bounds $\triangle_{d}$, $\triangle_{b}$ and $\triangle_{v}$ under different shard sizes. By default, we will conduct experiments based on these values.
TABLE III: Generic comparison of SP-Chain with several existing blockchain sharding systems.
System Performance Security
Throughput (tx/s) Latency (s) Unbiased randomness Resist malice during High efficiency
cross-shard tx under malice
Monoxide (1800 nodes) 600 15 Do not require $\surd$
RepChain (1800 nodes) 5628 58.2
RapidChain (1800 nodes) 4220 8.5 $\surd$
RapidChain (4000 nodes) 7380 8.8
SP-Chain (1800 nodes) 6650 6.4 $\surd$ $\surd$ $\surd$
SP-Chain (4000 nodes) 10305 7.6
7.2 System Throughput
We evaluate the throughput and scalability of SP-Chain under different scales. Fig. 5 shows the comparison results, where the purple bar is SP-Chain, and the yellow bar is Harmony. The figure shows that the throughput of SP-Chain can reach up to 2.8×2.8$\times$2.8 × of Harmony. Under the network scale of 4,000 nodes, SP-Chain’s throughput is 2.6×2.6$\times$2.6 × of Harmony (10,305 vs 3,901). Moreover, according to the Table III, SP-Chain’s throughput is better than several other existing works. Furthermore, Fig. 5 shows that the throughput of SP-Chain increases linearly with the expansion of the network scale. The above results indicate the superior throughput and scalability of SP-Chain.
7.3 Transaction Latency
In this section, we evaluate the transaction confirmation delay of SP-Chain. Figure 6 shows the comparison results. The transaction confirmation delays in SP-Chain are all less than half of that in Harmony. Under the network scale of 4,000 nodes, the delay of SP-Chain is only 0.48 (7.6s vs 15.7s) of Harmony. As the number of shards increases, the confirmation latency in both SP-Chain and Harmony increases as well. This is mainly because that as the number of shards increases, the number of nodes inside each shard also increases, so the intra-shard consensus takes longer time to finish. Furthermore, according to the Table III, the transaction delay of SP-Chain is also the lowest in those existing works.
7.4 System Decomposition

### v2

@@ VI-A Epoch Security
VI-B Analysis of Randomness Generation and Leader Rotation
We formalize the security of our chain-based distributed randomness scheme for leader rotation. In particular, we prove that the scheme’s output is unbiasable (fair) and unpredictable under standard cryptographic assumptions, meaning no adversary can manipulate or guess the leader selection beyond a negligible probability. We also note that the randomness is publicly verifiable and the generation process is scalable (incurring minimal overhead) [41].
Definition 1 (Unbiasability/Fairness).
[…]
Definition 2 (Unpredictability).
A scheme’s output is unpredictable if, prior to its generation, no computationally bounded adversary can guess the output with more than negligible advantage over random chance. For leader rotation, this means that until the randomness is revealed (by the confirmation of the current block), the identity of the next leader cannot be predicted except with probability close to $1/k$.
Theorem 1 (Fairness of Leader Selection).
Under the assumption of an honest majority in each shard (at most $f<k/2$ Byzantine nodes in a shard of size $k$) and an existentially unforgeable digital signature scheme, the leader selection in each slot of SP-Chain is unbiased. In other words, the randomness derived from the previous block’s signature yields a uniform random leader except with negligible probability.
Proof.
The leader for each new slot is determined by applying a public random oracle $H(\cdot)$ (modeled as a cryptographic hash function) to the prior confirmed block’s signature (and slot number). Because the slot number and block contents are fixed once a block is finalized, the input to $H$ is fixed and outside an adversary’s control at the time of leader election. A malicious leader cannot bias this process: given a unique identity and key pair per node (enforced by Sybil-resistant identity issuance and periodic reconfiguration of shard membership), a leader can produce at most one valid signature on a given block. Any attempt to produce an alternate seed (e.g., by forging another node’s signature or using an illegitimate identity) would require breaking the signature scheme’s unforgeability or introducing a fake identity, both of which are assumed infeasible (negligible probability). Furthermore, since $H$ is random-oracle unpredictable, the output is uniformly distributed over the space of possible leader identifiers. No biased influence can be exerted by the adversary beyond choosing not to sign or withholding the block (which would simply forfeit their turn). Thus, each eligible node has equal chance to be chosen by the hash, and no adversary (up to $f$ nodes) can skew this distribution except with negligible probability (e.g., by controlling a proportion of identities, which is bounded by $f/k$ for each shard). This establishes unbiasability of the leader rotation mechanism. ∎
Theorem 2 (Unpredictability of Leader Selection).
The signature-based randomness used for leader rotation is unpredictable until the moment of its use. Formally, even a Byzantine adversary (with up to $f$ corrupted nodes per shard) cannot obtain any information about the next slot’s leader prior to the decision of the lastet confirmed block, except with negligible probability.
Proof.
[…]
Verifiability and Scalability. The randomness generation is publicly verifiable – all inputs (the previous block hash/signature, public slot number, and the hash function) are publicly known, so any node or third party can independently recompute and verify the leader selection outcome. Moreover, the scheme is scalable: deriving randomness from existing blockchain data avoids any extra communication rounds. Each node simply performs a local hash computation to determine the leader, incurring $O(1)$ communication overhead (in contrast to heavier protocols like VSS-based beacons). This means our leader rotation mechanism adds virtually no networking cost while providing cryptographically strong randomness.
VI-C Analysis of Intra-Shard Consensus
We now rigorously define and prove the safety and liveness properties of the intra-shard consensus protocol (the pipelined two-phase concurrent voting combined with periodic leader rotation). We assume a synchronous network model with known maximum message delay $\Delta$, and that each shard has an honest majority ($f<k/2$ Byzantine nodes out of $k$). Under these conditions, we show that the consensus protocol maintains consistency (all honest nodes agree on the same shard state) and makes progress (new blocks eventually commit).
Definition 3 (Safety – Consistency).
A consensus protocol satisfies safety (or consistency) if no two honest nodes ever decide different blocks in the same round/slot. In a blockchain shard, this means that at the end of each slot, all honest shard members have committed the same block (or know that no block was committed in that slot). Equivalently, there are no "forks" within a shard’s ledger visible to honest participants.
Definition 4 (Liveness).
Liveness means that the protocol continues to make progress despite adversarial interference. For the shard consensus, liveness requires that any valid transaction submitted will eventually be included in a committed block. More concretely, in every epoch (or within a bounded number of slots), if the current leader is malicious or fails, the protocol will still eventually commit a new block under a future leader. This ensures that the shard’s chain grows and honest transactions are not starved indefinitely.
Theorem 3 (Safety and Liveness of Intra-Shard Consensus).
Under the synchronous network and honest-majority assumptions, the two-phase concurrent voting protocol with automatic leader rotation achieves both safety and liveness: (i) Safety: All honest nodes in a shard agree on the same block each slot (no conflicting commits occur). (ii) Liveness: The consensus will successfully commit new blocks in a timely manner, even in the presence of malicious leaders, as long as at least one leader in a succession of slots is honest (which is guaranteed by the rotation and majority assumption).
Proof.
Safety: We first consider a single slot with a designated leader. When an honest leader proposes a block, all honest replicas will receive the block digest (hash) within time $\Delta$ (by synchrony). Due to the two-phase voting design, an honest node waits a fixed pre-vote period (e.g., $\Delta_{b}$ as specified) after receiving the proposal before casting its vote, to allow slower nodes to catch up. In the worst case, the fastest honest node receives the block at time $t_{0}$ and the slowest at $t_{0}+\Delta$ (with $\Delta\leq\Delta_{d}$ in the parameters of our protocol). By the time the slowest honest node begins voting (after waiting $\Delta_{b}$), the fastest honest node has already broadcast its vote. All honest votes thus propagate and are received by every honest node within an additional $\Delta$ (bounded by $\Delta_{v}$). By the end of the allotted waiting time (on the order of $\Delta_{d}+\Delta_{b}+\Delta_{v}$ for the protocol parameters), every honest node has collected votes from all other honest nodes. A quorum of >$k/2$ consistent votes for the proposed block is reached, and all honest nodes will commit that block. No conflicting block can gather a quorum because either (a) there was a single proposal that all voted on, or (b) if a Byzantine leader tried to send divergent proposals (equivocation), honest nodes would detect the mismatch in block digests and immediately truncate the waiting period to vote against the faulty leader’s block. In an equivocation scenario, honest votes will not favor two different blocks: at worst, honest nodes reject the leader’s proposals altogether, resulting in no block for that slot rather than a fork. Additionally, if a malicious node attempts to cast two different votes, this double-voting is identifiable (each vote is signed by the node); honest nodes discard conflicting or invalid votes, so only one vote from each honest node counts. With an honest majority, no conflicting block can accumulate enough honest votes to be considered decided. Therefore, it is impossible for two different blocks to both reach commit threshold in the same shard, preserving safety.
Liveness: The automatic leader rotation ensures that even if the current slot’s leader is Byzantine (e.g., fails to send a proposal or sends an invalid one), the protocol will not deadlock. If a leader equivocates or sends malformed data, honest nodes detect it (by invalid signatures or inconsistent digests) and treat that block as failed; the slot will pass with no commit. Thanks to frequent rotation, a new leader will be selected in the next slot. Given the honest majority, with overwhelming probability an honest leader will be chosen after at most a few rotations (indeed, the probability of continuously picking malicious leaders for many consecutive slots is negligible). When an honest leader eventually takes charge, they will propose a valid block and the above safety argument shows all honest nodes will commit it. Even in the case of a silence attack (a malicious leader simply does nothing), all honest nodes will detect the absence of a valid proposal by the end of the slot duration and move on to the next leader. Thus, a faulty leader may delay progress for at most one slot, but cannot prevent an honest leader in a subsequent slot from committing a block. As long as at least one out of any consecutive set of leaders is honest (which is guaranteed given $f<k/2$), the shard will continue to append new blocks. This guarantees liveness: every valid transaction will eventually be processed when it is proposed under an honest leader (or a malicious leader’s turn passes without effect, and an honest leader later includes the transaction).
In summary, the intra-shard consensus protocol maintains one consistent chain per shard (safety) and will always produce new blocks (liveness) under the stated assumptions. Byzantine behavior such as equivocation, duplicate voting, or silence are tolerated by design: honest nodes either circumvent the misbehavior within the same slot (e.g., by accelerating the voting upon detecting equivocation) or move to the next slot’s leader, ensuring the protocol never violates safety and eventually commits all pending transactions. ∎
VI-D Analysis of Cross-Shard Transactions
Finally, we formalize the security guarantees of our cross-shard transaction processing mechanism and discuss its overhead. Cross-shard transactions in SP-Chain are designed to achieve atomicity (the transaction executes in all involved shards or not at all, eventually) and integrity (no tampering or inconsistency in transaction data across shards), even under Byzantine participants. We assume the cryptographic primitives are secure (hash functions are collision-resistant and signatures are unforgeable) and each shard has an honest majority as before.
Definition 5 (Atomicity of Cross-Shard Transactions).
A cross-shard transaction protocol provides eventual atomicity if for every transaction involving a source shard (origin of funds/assets) and a destination shard, it guarantees that either both of the following hold in the long run: the source shard’s state reflects the withdrawal (debit) and the destination shard’s state reflects the corresponding deposit (credit), or neither shard’s state reflects a completed transaction (no debit nor credit takes permanent effect). In simpler terms, funds or assets are neither lost nor duplicated: a cross-shard transfer will eventually complete in full, or be as if it never occurred.
Definition 6 (Integrity of Cross-Shard Transactions).
Integrity means that the content of a cross-shard transaction cannot be altered or forged during its transfer between shards. The destination shard will only accept the deposit if it exactly matches the transaction that was confirmed in the source shard. Any attempt by an adversary to modify the transaction amount, recipient, or any associated data, or to introduce a fake cross-shard transaction, must fail verification and be rejected. This property ensures the consistency of transaction semantics across shards.
Theorem 4 (Security of Cross-Shard Transaction Processing).
In SP-Chain’s cross-shard protocol, assuming honest majorities and secure cryptography, all cross-shard transactions satisfy atomicity and integrity. Specifically: (i) Atomicity: For any cross-shard transaction that debits an account in a source shard, the corresponding credit in the destination shard will eventually be committed (even if adversaries cause delays), so that the transaction’s net effect is all-or-nothing. (ii) Integrity: No adversary can tamper with the transaction data or execute a cross-shard transfer illegitimately; the deposit will only be accepted if it is authorized by a confirmed withdrawal in the source shard (as proven by cryptographic proofs).
Proof.
SP-Chain’s cross-shard execution follows a two-phase approach (inspired by atomic commit protocols and prior shard designs [44, 42]): the transaction’s withdrawal phase occurs in the source shard, then a deposit phase in the destination shard. Once the source shard commits the withdrawal (deducting the sender’s funds and logging the intent to transfer), a cryptographic proof of this event is relayed to the destination shard. This proof includes the Merkle inclusion proof of the transaction in the source shard’s block and the digital signatures attesting that block’s validity. The security properties follow from this design:
Atomicity: If the withdrawal sub-transaction in the source shard fails to commit (e.g., the source leader never includes it, or the source consensus does not finalize it), then naturally no funds are deducted and the cross-shard transfer has no effect (the transaction can be retried or aborted without inconsistency). If the withdrawal does commit in the source shard, the protocol guarantees that the deposit will eventually commit in the destination shard. Even a malicious leader in the destination shard cannot permanently stall the deposit: because leaders rotate, an honest leader in that destination shard will eventually be elected and will process the pending deposit. At worst, a Byzantine leader might delay including the deposit, but they cannot cause the source’s withdrawal to revert – the funds are already moved out of the source account and are effectively held awaiting deposit. Clients are aware of the status (they receive confirmation from the source shard) and can inform the destination shard or retry the transfer message if needed. Thus, with an eventually honest leader and persistent retransmission, the deposit will be executed. The net result is that after a finite delay, the destination account is credited. The transaction is never half-completed: there is no scenario where the source deducted funds but the destination never receives them – the worst-case outcome is a delay until an honest node finalizes the deposit. This meets the condition of eventual atomicity. (Likewise, it is impossible for a deposit to occur without its corresponding withdrawal: see integrity below.)
Integrity: The integrity is ensured by cross-shard verification using cryptographic proofs. When the destination shard receives the deposit request, it requires a proof that the corresponding withdrawal was confirmed in the source shard. This proof is the Merkle branch of the transaction in the source shard’s block, along with that block’s Merkle root and the source shard’s block header (which contains signatures from the source shard’s committee). Every node in the destination shard maintains an updated shard member table listing the public keys of all other shard committees (this table is refreshed each epoch during reconfiguration). Using this, the destination shard’s nodes can authenticate the source block’s signatures and thus trust the Merkle root. If a malicious actor alters the transaction (e.g., changing the amount or recipient in the deposit message), the Merkle proof will not match the altered data, and honest nodes in the destination shard will reject it. Similarly, an adversary cannot forge a valid proof for a transaction that never occurred in the source shard: without control of the majority of the source committee (which we assume they lack), they cannot produce a block header with a valid collective signature for a fabricated transaction. Any mismatch in signatures or Merkle roots is detected and the deposit is refused. Therefore, the only way a deposit is accepted is if it exactly corresponds to a legitimately confirmed withdrawal in the source shard. This guarantees that the state updates in the two shards are consistent and no cross-shard transaction can be modified or injected by Byzantine participants. ∎
@@ VI-D Analysis of Cross-Shard Transactions
VII Implementation and Evaluation
VII-A Experimental Setup
We implement SP-Chain based on Harmony [42] in Go language with 5,000+ lines of code. We implement BLS aggregated signature [3] in the prototype to reduce the signature size for better performance. For the baselines, we choose several notable existing blockchain sharding systems for performance comparison, including Monoxide [44], OmniLedger [21], and RapidChain [48]. Actually, the main protocols in SP-Chain can be easily applied to most existing sharding systems to improve system performance.
[…]
@@ VII-A Experimental Setup
Choice of Shard Size. We first determine the number of nodes per shard in SP-Chain based on Equation 2. The rule for selecting the shard size is: the failure probability of each shard should be less than $2^{-20}\approx 9\cdot 10^{-7}$ (time-to-failure of more than 4580 years for one-day epoch) [8, 48]. Table I shows the choice of shard size under different shard numbers and the corresponding failure probability. Results show that our choice of shard size makes the probability of failure less than $9\cdot 10^{-7}$ at any scale, ensuring the security of the system. The following experiments will be conducted based on the shard size determined by Table I.
@@
Choice of $\triangle$. The broadcast upper bound latency of digest, block and vote ($\triangle_{d}$, $\triangle_{b}$ and $\triangle_{v}$) are the 3 most important parameters in our system. To determine them, we evaluate the actual time spent in broadcasting digests, blocks, and votes in the system, and show the experimental results in Figure 5. Specifically, the bottom/top of each bar is the minimum/maximum delay in the broadcast, and the horizontal line connected to each bar is the average delay. Based on the measured actual broadcast delay, we set the broadcast delay upper bound as $delay_{max}+\sigma$. Where $delay_{max}$ is the actual maximum broadcast delay, $\sigma$ is the standard deviation obtained by fitting the broadcast delay to a normal distribution. Table II shows the calculated delay upper bounds $\triangle_{d}$, $\triangle_{b}$ and $\triangle_{v}$ under different shard sizes. By default, we will conduct experiments based on these values.
VII-B System Throughput
We evaluate the system throughput and scalability of SP-Chain under varying network sizes from 680 up to 4,000 nodes, in direct comparison with Monoxide, OmniLedger, and RapidChain. Figure 6 presents the throughput results across these scales. SP-Chain consistently achieves the highest throughput at every network size. For example, on a 680-node network SP-Chain sustains approximately 3,500 TPS, which already significantly exceeds the throughput of RapidChain (1,900 TPS), OmniLedger (1,200 TPS), and Monoxide (400 TPS). As the network size grows, this gap widens. Under the largest tested scale of 4,000 nodes, SP-Chain processes 10,305 TPS, whereas RapidChain and OmniLedger reach only about 7,400 TPS and 3,900 TPS respectively, and Monoxide remains 2,600 TPS. In other words, at 4,000 nodes SP-Chain achieves roughly 1.4×1.4$\times$1.4 × the throughput of RapidChain, about 2.6×2.6$\times$2.6 × that of OmniLedger, and around 4×4$\times$4 × that of Monoxide. Furthermore, SP-Chain’s throughput increases almost linearly with the expansion of network size, demonstrating superior scalability.
VII-C Transaction Latency
We now compare the average transaction confirmation latency of SP-Chain against baselines as the network size increases. Figure 7 shows the confirmation delay results for network sizes from 680 to 4,000 nodes. Across all scales, SP-Chain achieves substantially lower latency than the other protocols. At the 680-node scale, a transaction in SP-Chain is confirmed in roughly 6s, whereas RapidChain requires about 8s, OmniLedger about 13s, and Monoxide around 15s on average for confirmation. As the network grows, the confirmation latency of every protocol rises due to larger shard sizes and increased cross-shard coordination. SP-Chain’s latency, however, grows very modestly and remains the lowest among the systems. In a 4,000-node network, SP-Chain confirms transactions in about 7.6s, which is roughly half of that in OmniLedger and Monoxide, and about 15% lower than in RapidChain. These results demonstrate that SP-Chain sustains low confirmation latency even as the system scales, outperforming prior protocols on this metric. Combining throughput and latency, our evaluation shows that SP-Chain does not trade off latency for throughput – it delivers both higher throughput and faster confirmations than various baselines across the range of network sizes tested.
VII-D System Decomposition